How many years an investment would be four times its initial value at an interest rate of 11 percent.


1
Expert's answer
2021-04-18T19:28:43-0400

Assume the investment is X;

r=0.11

Use grow it using the following formula.

"X(1+r)^n=4X"

Divide both sides by X;

"\frac{X(1+r)^n}{X}=\frac{4X}{X}"


"(1+r)^n=4"

Substitute for r;

"(1.11)^n=4"


Introduce logarithmic

"n log 1.11=log 4"


"n= \frac{log 4}{log 1.11}=13.28 years"


Investment will be 4 times after a lapse of 13.28 years.